OpenStudy (anonymous):

Please help! Find the zeros of the function f(x) = x^3 + 2x^2 + x - 2

OpenStudy (campbell_st):

well I just graphed it... and it has 1 zero...

OpenStudy (campbell_st):

and given f(0) = -2 and f(1) = 2 the zero is between x = 0 and x = 1 so perhaps the half interval method or Newton's method could be used to approximate the zero...

OpenStudy (anonymous):

I've never heard of either of those methods. Hey can I do this: if I find the zeros for g(x) = x^3 + 2x^2 + x 0 = x(x^2 + 2x + 1) x = 0, x = -1 But then subtract those by -2 since g(x) is a vertical translation of f(x) So the zeros of f(x) are x = -2 and x = -3? But isn't there a way to solve this using algebra?

OpenStudy (anonymous):

Actually I graphed it and it says the zero is at 0.7 but I have no idea how to get that answer

OpenStudy (campbell_st):

ok.... well a really easy solution is to graph the curve... that will give you the zero... well the translation is down 2 so the zero of x =-1, as its a perfect square so only touches at x = -1, will now disappear... and the zero at x = 0 moves right....

OpenStudy (anonymous):

So there isn't a way to solve this using basic algebra or algebra 2?

OpenStudy (anonymous):

Like the way you'd solve another function that can be factored or something

OpenStudy (campbell_st):

well the problem is the curve doesn't have a rational root...

OpenStudy (campbell_st):

using the rational root theorem the possible rational roots are -1, 1, -2, and 2 none of those give f(x) = 0 so the root, is irrational...

OpenStudy (anonymous):

That makes sense

OpenStudy (campbell_st):

there is probably a solution, but I haven't come across it

OpenStudy (anonymous):

So I should estimate it as 0.7 because I can't solve it algebraically right?

OpenStudy (anonymous):

We haven't learned that integral calculus stuff you said

OpenStudy (campbell_st):

well I graphed it using https://www.desmos.com/calculator and it gave me x = 0.696

OpenStudy (campbell_st):

newton's method is differential calculus... its used to approximate the roots by using iterations.

OpenStudy (anonymous):

fun stuff I"ll have to learn later. Well thanks for helping me with this problem :]

OpenStudy (campbell_st):

good luck